OpenStudy (anonymous):

What is the slope of the line x = 4

OpenStudy (anonymous):

slope of infinity its a straight line

OpenStudy (anonymous):

the slope =(y2-y1)/(x2-x1)= since x is a constant (y2-y1)/0 = infinity

OpenStudy (anonymous):

Thanks